HELP ME..
juan and peter both leave a shopping center at the same time going in opposite directions. Juan is on his bike and travels 3mps faster than peter who is on his skateboard. after 1.5 hours they are 19.5 miles apart. How fast does peter travel?
answer is.... 5mph
I NEED THE SOLUTION :(

Answer by checkley77(12569)

(X+X+3)1.5=9
(2X+3)1.5=9
3X+4.5=9
3X=9-4.5
3X=4.5
X=4.5/3
X=1.5 MPH FOR SKATEBOARDER.
1.4+3=4.5 MPH FOR THE BIKER.
PROOF:
(1.5+1.5+3)1.5=9
6*1.5=9
9=9